INAF LBC+LUCIFER queue observing, 2010 November 30 - December 01

Observer: K.Boutsia, M. Fumana,V. Testa
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC - LUCIFER

Summary

Clear Sky at the beginning, some cirrus at the horizon
20min lost due to software problems with dofpia - it could not find a library since it was pointing to an older build that had been deleted
Possible clouds during the night - but in general good seeing (0.8'' -1'' for LBC and less than 0.6'' for LUCIFER) ????CHECK
Attention - Imaging Flats taken with LUCIFER at dawn have been mistakenly called luci_20101011_FLat_... Difficult to find in archive. We copied them to external disk and renamed them

Data Summary

Project Objects Filter ExpTime Notes
Clementini AndXXI? - AndXIX? B-bessel, V-bessel 105min + 105min seeing <1.0' - some elongation problems at the beginning
Turatto SN2009kr J, H, Ks 640 s + 896 s + 896 s  
Turatto SN2010jl J, H, Ks 12 s + 12 s + 24 s guiding probe covers 1.4 of field during 3rd offset
Turatto SN2010jl 200HK 2000sec Tell+ARC+FLAT
Cresci HDF 210zJHK+H 1hr ARC+FLAT - NO Tell - use the one in Dec02

Details

Begin night with LBC

12:35 - Skyflat Test LBC - B(blue) - V(red) ok

12:41 - SkyFlat? B-V
  • 1 sec B - V
  • 2 sec B - V (increased to 4 during the OB)
  • PA 180 1.5 sec B - 0.5 sec V
  • PA 180 - 6sec B + 6sec V

12:53 - Slew to pointing star ACT0219
  • Pointing correction OK + mirror adjust OK
  • problems with setting the camera name - David fixed it...

01:10 - slew to PHOT_STD - SA113 dofpia,
  • manually give Z11 = +1000
  • problems with the dofpia - does not find the file libTCS.so -call Olga and call Norm -
  • chris moved some files around - 01:31 we start dofpia again
  • dofpia,
  • Z7 = -2000 , Z8 =-1000, Z4=+10000 (only Red)
  • keep trying to focus manually...

01:49 - slew to Clementini - AndXXI? -dofpia
  • hit limits on the red mirror

02:04 - observing target Clementini - M31_AndXXI_420s_1
  • Blue images slightly elongated. Seeing good (0.8 V, 1.0 B)
  • Sky clear, thin clouds at horizon SW
  • Red Master

02:32 - M31_AndXIX_420s_1
  • dofpia

02:38 - observing target Clementini - M31_AndXIX_420s_1
  • very high elevation - not co-pointing correctly because blue mirror is hitting limits
  • images in both sides look elongated
  • it crossed the meridian during the last exposure of the OB
  • OB probably out of specs

03:02 - slew to AndXXI - dofpia

03:09 - observing target Clementini - M31_AndXXI_420s_2

03:33 - slew to AndXIX - dofpia

03:39 - observing target Clementini - M31_AndXIX_420s_2

04:03 - slew to AndXXI - dofpia

04:08 - observing target Clementini - M31_AndXXI_420s_3

04:19 - slew to AndXIX - correct mirror - dofpia

04:46 - observing target Clementini - M31_AndXIX_420s_3

05:10 - slew to AndXXI - dofpia
  • after 5th iteration blue still did not converge but it was very close
  • we send current values and start with science exposure

05:21 - observing target Clementini - M31_AndXXI_420s_4

05:47 - slew to AndXIX - dofpia

05:53 - observing target Clementini - M31_AndXIX_420s_4

06:18 - slew to AndXXI - dofpia

06:24 - observing target Clementini - M31_AndXXI_420s_5
  • airmass: reached 1.53

06:48 - slew to AndXIX - dofpia

06:52 - observing target Clementini - M31_AndXIX_420s_5
  • airmass: 1.5 to 1.6

07:15 - SWITCH TO LUCIFER
  • Pointing Correction and collimation done by TO
  • seeing is ~ 0.6 on the guider

08:07 - Slew to SN2009kr image acquisition (Turatto) 0003-0004
  • we changed camera from N1.8 to N3.75
  • also offset from relative to absolute ...
  • sky reasonably clear, maybe thin cyrrhus here and there

08:15 - Science imaging scripts for SN2009kr
  • Imaging H 0005-0020
  • Imaging J 0021-0028
  • Imaging Ks 0029-0044 (started at 08:56)

09:34 - Slew to SN2010jl image acquisition (Turatto) 0045-0046
  • scripts had to be slightly modified

09:41 - Science imaging scripts for SN2010jl
  • Imaging H 0047-0050
  • Imaging J 0051-0054
  • Imaging Ks 0055-0058
  • in the 3rd offset the probe of the guiding star covers 1/4 of the field

09:50 - Spec acquisition script for SN2010jl 0059-0065

10:18 - Spec science scripts SN2010jl 0066-0073

10:55 - Calibration Unit IN
  • ARC 0074-0075
  • FLAT 0076-0111 check

11:15 - Slew to Telluric for SN2010jl spectra
  • acquisition 0112-0117

11:27 - Science script for the Telluric 0118-0121

11:32 - Calibration Unit IN *FLAT 0122-0130

11:48 - Slew to HDF (Cresci)
  • MOS aqcuisition script 0131-0135

12:08 - Science MOS script for HDF in the H band 0136 -0147
  • NO TELLURIC because sky too bright

13:17 - Calibrationi Unit IN
  • FLAT H 0148-0153
  • ARC H 0154-0157

13:25 - Move to Flat position
  • used script to make the flats and they were saved as luci_20101201_Flat_????
  • Flats in J: 001-010 still dark, 011-020 low (1000), 031-040 ( 2500-3500), 051-060 high(7500-11000)
  • Flats in H: 021-030 low (1000-1500), 041-050 ( 2800-4500), 61-70 high(9000-14000)
  • Flats in Ks: 071-100 (2000-4800)

13:50 - END OF NIGHT
  • Darks sciprt

-- VincenzoTesta - 01 Dec 2010---
Topic revision: r9 - 15 Mar 2012, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback